6n+4n-n+5=23
36-4c-3c=22
28=10a-5a-2
3/8+p/8-31/8
4g/14-3/7-g/14=3/14
5/11=-3b/11+8b/22

Answer by checkley77(12844)

6n+4n-n+5=23
9n=23-5
9n=18
n=18/9
n=2 answer.
---------------------
36-4c-3c=22
-7c=22-36
-7c=-14
c=-14/-7
c=2 answer.
----------------------
I've solved the first two to show you what to do so you should be able to finish the assignment.